The problem is: Two planes are 6000 miles apart, and their speeds differ by 200 miles per hour. They travel toward each other and meet in 5 hours. Find the speed of the slower plane.
Answer by ankor@dixie-net.com(12699)

Two planes are 6000 miles apart, and their speeds differ by 200 miles per hour.
They travel toward each other and meet in 5 hours.
Find the speed of the slower plane.
:
let s = speed of the slower speed
then
(s+200) = speed of the faster
:
Write a distance equation: dist = time * speed
:
slow plane dist + fast plane dist = 6000mi
5s + 5(s+200) = 6000
5s + 5s + 1000 = 6000
10s = 6000 - 1000
10s = 5000
s = 500 mph is the slow speed
;
:
Check solution (fast speed plane = 700 mph
5(500) + 5(700) =
2500 + 3500 = 6000